OpenStudy (anonymous):

1. If I know the lengths of two sides of a right triangle, how do I find the third? 2. Could I find the two missing side lengths of a right triangle if I only know one side length and one angle measure (other than the 90 degree angle)? 3. Could I find the two missing angle measures if I know some of the side lengths of a right triangle? 4. What makes a triangle a “special” right triangle? How can special right triangles help me find side lengths?

OpenStudy (anonymous):

•Answer the question in complete sentences, as if you were explaining it to your friend. •Create a unique teaching example to demonstrate what you explained in words. Work through this example step by step. • Create a unique practice example. Provide the problem and the answer. Because you'd like your friend to show their understanding of what you taught, make the practice example different from the teaching example. For example, if the teaching example for #1 was missing the hypotenuse, have the practice example missing one of the legs. Or, if the teaching example for #4 used the sin ratio, have the practice example use the cosine or tangent ratio. Note: An example or practice problem is considered unique when it has not been copied from another source, including the lesson, practice, assessments, or Internet.

OpenStudy (anonymous):

i need help dont understand geometry that much

OpenStudy (anonymous):

yes you can find it using trigonometry if one angle is given. and if two sides are given use pythagoras thm.

OpenStudy (anonymous):

is that number one and can u plzz help me with this i no hardly anything about trigonometry and all

OpenStudy (anonymous):

let a be any angle given in triangle then side opposite to a is called as perpendicular(p) , side opposite to 90 deg angle is called as hypotenuse(h) and remaining side is base(b) tan a =(p/b) sin a =(p/h) cos a =(b/h)

OpenStudy (anonymous):

thank you soo much are you doing these in order?? it seems like it to me..